# NeRFmentation: NeRF-based Augmentation for Monocular Depth Estimation

### v1

@@ 2 Related work
AdaBins 0.561 0.851 0.933 0.224 2.201 8.027 0.328
+Reconstructed 0.658 0.879 0.944 0.204 2.056 7.397 0.292
+Interpolated 0.681 0.874 0.940 0.200 2.156 7.389 0.286
+Angled ±$\pm 3^{\circ}$ 0.632 0.855 0.929 0.215 2.237 7.835 0.313
DepthFormer 0.545 0.885 0.958 0.230 1.920 7.776 0.313
+Reconstructed 0.607 0.907 0.969 0.199 1.723 7.730 0.275
+Interpolated 0.669 0.926 0.975 0.184 1.542 6.755 0.252
+Angled ±$\pm 3^{\circ}$ 0.602 0.909 0.970 0.199 1.711 7.286 0.274
BinsFormer 0.571 0.901 0.966 0.208 1.782 7.547 0.289
+Reconstructed 0.702 0.924 0.973 0.183 1.598 6.714 0.249
+Interpolated 0.707 0.928 0.975 0.181 1.556 6.563 0.244
+Angled ±$\pm 3^{\circ}$ 0.680 0.925 0.975 0.185 1.573 6.866 0.251
Table 1: AdaBins, DepthFormer, and BinsFormer trained on augmented KITTI dataset evaluated on the Waymo dataset. We showcase the performance of AdaBins [6], DepthFormer [16], and BinsFormer [15] models trained on the KITTI Eigen [9, 12] train split and compare it to the NeRFmented (Ours) versions of the same models evaluated against the unseen Waymo Open Dataset. Our method greatly increases the robustness to unseen data, as demonstrated by the increased performance of NeRFmented models on the Waymo dataset in comparison to the baseline models which were trained without introducing novel views. The best results are in bold dark green. Second best results are underlined light green.
Data augmentation.
Data augmentation is a well-established regularization technique in Machine Learning that generally enables the training of increasingly complex models by applying software-based label-preserving transforms to scarce data. These augmentations for Convolutional Neural Networks (CNNs) and derived models started off with simple transformations such as random translations, rotations, or flips [32, 13]. Recent advancements in generative modeling have enabled the creation of photo-realistic synthetic data through increasingly sophisticated data augmentation techniques. These methods are based on Generative Adversarial Networks (GANs) [27, 5] or diffusion models [2, 1]. Sandfort et al. employ CycleGAN [43], a GAN that enforces cycle-consistency, to augment images for medical CT segmentation tasks [27] and successfully improve the performance even for out-of-distribution samples. While Rajpal et al. show in [21] that it is possible to improve monocular depth estimation performance on real-world data using a fully synthetic custom dataset, we aim to show that real-world datasets can be extended synthetically using NeRFs to improve the performance. Yen-Chen et al. [40] use NeRFs for robotic action planning by building a 3D scene representation and subsequently using novel view synthesis of orthographic images. The authors are able to show that compared to images produced by physical cameras, these synthetic orthographic images allow for better pixel-wise action values correlated with picking and placing success of objects by the robot.
Mancini et al. [17] argue for the need for generalization of MDE networks for autonomous mobile robots, to “guarantee robust operation in real-world scenarios”. To achieve this, they train on two synthetic datasets, one urban environment and one forest environment. They then test the models trained on the synthetic datasets on KITTI, where they achieve results comparable to Eigen et al. [9]. They also test on two datasets they collected themselves to test generalization, one is of countryside roads, and one is of a forest. They find that the network trained on both urban and forest environments performs better than networks trained only on one of the scenarios.
Bauer et al. [4] use novel view synthesis to improve depth prediction. They use three blocks: one network to predict depth, a block that transforms views from one direction to another, by using the available depth data, and one network which reconstructs the image obtained from the previous block, since it is imperfect. They manage to beat the state of the art in one metric on the KITTI dataset.
3 Methods
We propose a dataset augmentation technique for Monocular Depth Estimation which is, to the best of our knowledge, novel. We leverage the fact that many real-world image datasets have a sequence-like structure, meaning that the same environment is captured from many slightly different viewpoints. Furthermore, pose data for each frame is often either provided or can be inferred through structure-from-motion, for example using COLMAP [29, 28]. Using this data, we train NeRFs to encode 3D scenes and subsequently render novel RGB-D images from unobserved poses.
We apply our proposed augmentation method to the KITTI Eigen training split [12, 9] and create a ”NeRFmented” version of it to show the diversity of scenes to which our method can be applied. We train the off-the-shelf MDE models AdaBins [6], DepthFormer [16], and BinsFormer [15] on the NeRFmented dataset and evaluate them on the original dataset against their vanilla-trained model to show the transferability between different architectures. Finally, we evaluate the NeRFmented models vs the vanilla-trained models on a different dataset that the models have not seen during training to show the added robustness and diversity that our method introduces. The ablation dataset is the Waymo Open Dataset [35].
3.1 Proposed pipeline
For each scene in the source dataset, we train a separate NeRF using RGB-D images and their corresponding poses to encode a 3D representation. To ensure a high reconstruction quality for each NeRF, we withhold a small part of the input data from the training. Once the training has concluded, we use the withheld data as a small validation split to potentially filter out scenes with subpar image- or depth reconstruction quality. We use the retained NeRFs to subsequently render RGB-D images from novel viewpoints, which are typically obtained by perturbing the original poses either through translation or rotation. Finally, we combine the RGB-D images from our novel views with the source dataset to form an augmented dataset and train the downstream monocular depth estimation network on it. The overall pipeline is depicted in Figure 2.
3.2 Scene representation model
For the NeRFs of our pipeline we use the off-the-shelf Nerfstudio framework [37], which provides convenient access to various NeRF extensions. Because our method requires high-quality depth reconstruction and image reconstruction, we base our model on a combination of the depth-Nerfacto and Nerfacto-huge architectures, which were both introduced by Tancik et al. in [37]. This way, we benefit from depth-supervision that is included in the depth-Nerfacto training while increasing the learning capacity of the network by replacing the MLP of depth-Nerfacto with the larger MLP of Nerfacto-huge. Figure 3 shows the qualitative performance gain of our depth-Nerfacto-huge compared to the vanilla depth-Nerfacto. We implement a filtering step which eliminates NeRFs that fail to meet a specified dataset-dependent reconstruction quality standard. To achieve this, we utilize the trained NeRF for each scene to render the RGB-D images from the poses of the small withheld validation split and compare them to the corresponding ground-truth RGB-D images. This involves computing the absolute relative error to assess the depth reconstruction quality and the LPIPS score [41] to evaluate the image reconstruction quality. We define dataset-specific thresholds for both metrics and any NeRF that falls below these thresholds is discarded.
RGB
Refer to caption Refer to caption Refer to caption Refer to caption
AdaBins [6]
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFmented AdaBins [6]
Refer to caption Refer to caption Refer to caption Refer to caption
DepthFormer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFmented DepthFormer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative Results on the Waymo [36] dataset. We show the performance of NeRFmentation (Ours) compared to vanilla-trained AdaBins [6] and DepthFormer [16] leading to improved predictions by using NeRF augmented data for introducing additional viewpoints in the training images. Both the NeRFmented models and vanilla-trained models have been trained on the KITTI Eigen [9, 12] train split and evaluated on the unseen Waymo Open Dataset. Color scale: 0 (purple) to 80 meters (yellow).
@@ 3.2 Scene representation model
4 Experiments
4.1 Datasets
KITTI [12].
The KITTI dataset is an outdoor dataset designed for diverse autonomous driving scenarios, featuring stereo-image pairs and annotated LiDAR ground-truth depth data. The KITTI Eigen split [9] is a popular split of the dataset for monocular depth estimation with approximately 23,000 stereo-image pairs for training which are separated into 32 scenes for training and validation. The corresponding validation split consists of 888 stereo-image pairs. The test split contains 697 images from 29 different scenes.
Waymo Open Dataset [35].
The Waymo Open Dataset (Waymo) dataset is a large-scale outdoor dataset collected and developed by Waymo LLC for various autonomous driving tasks. The dataset incorporates data from five high-resolution pinhole cameras and five LiDARs and other sensors, which offers a wider range of camera viewpoints, denser depth maps, and a more thorough overview of the surroundings compared to datasets such as KITTI [12]. This enables more robust training of models used for autonomous driving tasks. To evaluate the robustness of our NeRFmented models, we test them against the Waymo test dataset. We choose to use only the subset of the test set coming from the front, front-left, and front-right facing cameras. We do this to avoid too large of a domain shift from the standard MDE datasets based on images taken by forward-facing cameras, such as KITTI.
4.2 Monocular Depth Estimation models
To show that our augmentation technique is model-agnostic, we show it in combination with three popular MDE models: AdaBins [6], DepthFormer [16], and BinsFormer [15]. We train and evaluate using the Monocular-Depth-Estimation-Toolbox [14], an open-source monocular depth estimation toolbox aimed at collecting several state-of-the-art MDE models into a single environment. Due to computational constraints, we can not use the original batch size used for DepthFormer and BinsFormer. Hence, we decrease the learning rate and increase the number of iterations accordingly.
For the evaluation of the MDE models, we use the same metrics as in [4]. For their definitions we refer to the supplementary material.
4.3 Training of NeRFs
We train our depth-Nerfacto-huge model on each scene of the KITTI Eigen [12, 9] train split and retain 10 % of the training data as a validation split. As we experimentally find that the reconstruction quality of NeRFs drastically reduces as the size of a scene increases, the varying size of the scenes in the KITTI dataset requires splitting them into sub-scenes. We do not want to decrease the scene size too much though, as this would increase the computational complexity and bring no additional gain when rendering novel views from challenging poses. In our case, we ensure that no pose in each sub-scene should be more than 50m away from the first pose in that sub-scene to solve the mentioned trade-off. For each sub-scene, we train a separate NeRF.
For each NeRF, we tune the weight of the depth loss and whether the poses are refined or not by running a grid search for 10,000 steps. We observe that when using NeRFs for 3D reconstruction, there is a strong trade-off between RGB quality and accuracy of the geometry, which in turn affects the accuracy of the depth maps. Since we require both to be of high quality, we use the small validation split that we withheld from the training and re-render the RGB-D images for each of their poses using all NeRFs from our grid search. Then, we select the NeRF that minimizes a trade-off validation measure
$T_{\text{RGBD}}=\sqrt{\alpha\cdot\text{Abs. Rel}^{2}+\mathrm{LPIPS}^{2}}$ (1)
where we use the absolute relative error of the reprojected depth maps as a quality measure of the geometry and the LPIPS score [41] as a quality measure of the RGB reconstruction. We set $\alpha=10$ for this experiment.
Finally, we continue training the best run of the hyperparameter tuning of each NeRF until we reach 30,000 iterations. Then, we apply the filtering step described in Section 3.2 using the same validation split where we discard all sub-scenes and their corresponding NeRF, where the NeRF did not achieve an LPIPS score of less than $0.22$ or an absolute relative error of less than $0.05$ on the 3D reconstructions. We choose the threshold for LPIPS empirically by observing which value produces minimally visually acceptable RGB images. The threshold for depth absolute relative error is chosen because it is an approximate lower bound of state-of-the-art MDE networks. We filter any scenes with a higher score so as not to introduce extraneous noise into the training dataset.
Figure 6 shows the images reconstructed by NeRFs alongside ground-truth images present in the KITTI dataset. Apart from a low LPIPS score, visual inspection of the reconstructed RGB images also suggests that they are of comparable quality to the original. Additionally, the reconstructed depth maps are dense, whereas the ground-truth ones are sparse. The low absolute relative error of the reconstructed depth maps indicates a high agreement between the original and reconstructed depth maps as well.
4.4 Novel view synthesis
We propose and evaluate three data augmentation strategies: (i) Reconstruction: Re-rendering the exact train poses used to train the NeRF. This completes the sparse depth from the source dataset. (ii) Interpolated: Creating novel poses by interpolating between pairs of training poses as these poses inject additional data while benefiting from good depth and RGB supervision. (iii) Angled: Rendering two additional views for each train pose, rotated ±$\pm 3^{\circ}$ horizontally as this injects the most diverse data while simulating a plausible motion performed by a driving car. For the KITTI Eigen train split [9, 12] this results in dataset sizes of (i) 29,836 (+28%), (ii) 29,744 (+28%), and (iii) 32,132 images (+39%).
The motivation behind these data augmentation strategies is to evaluate whether it is sufficient to use NeRFs as depth completion networks or whether certain novel views of training scenes can improve performance on evaluation datasets.
GT
Ours
GT
Ours
Figure 6: Qualitative NeRF reconstruction results on KITTI [12]. We show original images from the KITTI dataset and the corresponding images generated using the trained and filtered NeRFs for the same camera poses. It is evident that reconstructed RGB images are very close to their real counterparts. The NeRFs also complete the sparse ground truth depth maps.
4.5 Training on KITTI
We train the described off-the-shelf MDE networks on the original KITTI Eigen [12, 9] train split and on our NeRFmented version of it and compare the results based on the standard MDE metrics. Our models are trained from scratch using the same hyper-parameters as described in the corresponding papers to show the influence of our augmentation method compared to the baseline without additional factors contributing to the performance change.
Among the three MDE architectures trained with NeRFmented variants of the KITTI Eigen train split, in Table 2 we see that the DepthFormer models surpass the baseline model trained on the standard KITTI Eigen train split. The angled ±$\pm 3^{\circ}$ NeRFmentation yields an improvement in the majority of metrics we evaluate for, breaking even for the rest. The interpolated NeRFmentation is not far behind, yielding better results for 3/7 metrics while the reconstructed NeRFmentation is behind both, suggesting that the introduction of diverse poses through angle variation and interpolation introduces novel information that the MDE visibly benefits from.
[…]
$\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
AdaBins 0.964 0.995 0.999 0.058 0.190 2.360 0.088
+Reconstructed 0.961 0.995 0.999 0.060 0.204 2.434 0.092
+Interpolated 0.960 0.995 0.999 0.61 0.204 2.408 0.092
+Angled ±$\pm 3^{\circ}$ 0.961 0.995 0.999 0.061 0.204 2.412 0.093
DepthFormer 0.975 0.997 0.999 0.052 0.158 2.143 0.079
+Reconstructed 0.974 0.997 0.999 0.052 0.152 2.109 0.098
+Interpolated 0.976 0.997 0.999 0.052 0.150 2.072 0.079
+Angled ±$\pm 3^{\circ}$ 0.974 0.997 0.999 0.051 0.150 2.072 0.078
BinsFormer 0.974 0.997 0.999 0.052 0.151 2.098 0.079
+Reconstructed 0.974 0.997 0.999 0.054 0.155 2.104 0.081
+Interpolated 0.973 0.997 0.999 0.055 0.157 2.125 0.081
+Angled ±$\pm 3^{\circ}$ 0.973 0.997 0.999 0.055 0.159 2.164 0.082
Table 2: Comparison of performances on the KITTI Eigen [12, 9] test split. We showcase the performance of AdaBins [6], DepthFormer [16] and BinsFormer [15] models trained on the KITTI Eigen [12, 9] train split and compare it to the NeRFmented versions of the same models, using different novel view generation strategies. NeRFmented models stay very close to baseline performance. The best results are in bold dark green. The second best results are underlined light green.
4.6 Zero-shot dataset transfer to Waymo
@@
Evaluation on a different dataset provides a good insight into generalization capabilities. In this comparison, the models are strictly trained on KITTI and KITTI-based NeRFmented images. Table 1 shows that the NeRFmented runs deliver massive uplifts across the board, beating baseline performance in a large majority of metrics for every single augmentation strategy. Notably, interpolation outperforms reconstruction and angling most of the time, likely as it strikes a good balance between providing novel viewpoints while staying close to train data distribution.
The quantitative success is mirrored by the qualitative results displayed in Figures 4 and 5. As a result of the densification of the depth maps, the MDE architecture gains supervision in the upper part of the image. This effect can be seen in the increased clarity in the upper parts of the predicted depth maps. Additionally, an increase in fine detail prediction in street signs, stoplights, street lamps, and foliage is clearly visible. Car outlines become sharper and background building details are much improved as well.

### v2

@@ 4.4 Training of NeRFs
CAug. ✓ Adabins 0.561 0.851 0.933 0.224 2.201 8.027 0.328
 • FDA [39] 0.597 0.884 0.952 0.210 1.906 7.467 0.297
 • Style Transfer [1] 0.618 0.895 0.957 0.205 1.836 7.207 0.287
 • NeRFmentation 0.666 0.866 0.937 0.209 2.317 7.667 0.293
CAug. ✗ Adabins 0.543 0.854 0.938 0.228 2.248 8.562 0.330
 • FDA [39] 0.584 0.861 0.938 0.226 2.180 7.983 0.320
 • Style Transfer [1] 0.591 0.847 0.927 0.225 2.300 8.215 0.329
 • NeRFmentation 0.629 0.873 0.946 0.220 2.189 7.770 0.295
CAug. ✓ DepthFormer 0.545 0.885 0.958 0.230 1.920 7.776 0.313
 • FDA [39] 0.479 0.891 0.964 0.225 1.895 7.793 0.308
 • Style Transfer [1] 0.470 0.880 0.958 0.230 1.931 7.867 0.319
 • NeRFmentation 0.724 0.926 0.975 0.170 1.489 6.777 0.238
CAug. ✗ DepthFormer 0.479 0.882 0.961 0.232 1.987 8.006 0.313
 • FDA [39] 0.445 0.873 0.960 0.238 2.052 8.209 0.322
 • Style Transfer [1] 0.464 0.864 0.954 0.236 2.155 8.544 0.328
 • NeRFmentation 0.764 0.937 0.979 0.160 1.355 6.365 0.222
CAug. ✓ BinsFormer 0.571 0.901 0.966 0.208 1.782 7.547 0.289
 • FDA [39] 0.599 0.914 0.971 0.201 1.676 7.209 0.276
 • Style Transfer [1] 0.545 0.891 0.961 0.214 1.837 7.718 0.301
 • NeRFmentation 0.744 0.928 0.975 0.168 1.527 6.617 0.231
CAug. ✗ BinsFormer 0.615 0.905 0.968 0.202 1.746 7.505 0.279
 • FDA [39] 0.600 0.903 0.967 0.206 1.794 7.601 0.284
 • Style Transfer [1] 0.571 0.894 0.963 0.210 1.840 7.882 0.294
 • NeRFmentation 0.770 0.938 0.978 0.166 1.416 6.181 0.220
Fig. 4 shows the images reconstructed by NeRFs alongside ground-truth images present in the KITTI dataset. Apart from a low LPIPS score, visual inspection of the reconstructed RGB images also suggests that they are of comparable quality to the original. Additionally, the reconstructed depth maps are dense, whereas the ground-truth ones are sparse. The low absolute relative error of the reconstructed depth maps indicates a high agreement between the original and reconstructed depth maps as well.
Table 2: Ablation of NeRFmentation strategies for models trained on the augmented KITTI dataset evaluated on the Waymo dataset and the KITTI Eigen [10,14] test split. We showcase the performance of AdaBins [5], DepthFormer [17], and BinsFormer [18] models trained on the KITTI Eigen [8, 12] train split and compare NeRFmentation strategies. No classic augmentation is used. The best NeRFmentation strategy to use depends on the chosen depth architecture. The best results are in bold. The second best results are underlined.
Dataset Waymo KITTI
[…]
Adabins  • NeRFm. Reconstruction 0.667 0.893 0.956 0.201 1.932 7.335 0.276 0.958 0.994 0.999 0.062 0.210 2.506 0.094
 • NeRFm. Interpolation 0.680 0.875 0.944 0.197 2.024 7.352 0.281 0.957 0.995 0.999 0.061 0.213 2.524 0.094
 • NeRFm. Angled ±$\pm 3^{\circ}$ 0.669 0.886 0.953 0.204 2.040 7.470 0.280 0.958 0.994 0.999 0.062 0.213 2.513 0.095
 • NeRFm. Horiz. ±$\pm 30\,cm$ 0.684 0.884 0.952 0.202 2.039 7.593 0.279 0.955 0.993 0.999 0.065 0.234 2.595 0.098
 • NeRFm. Vert. ±$\pm 30\,cm$ 0.629 0.873 0.946 0.220 2.189 7.770 0.295 0.958 0.994 0.999 0.062 0.215 2.534 0.095
DepthFormer  • NeRFm. Reconstruction 0.735 0.932 0.977 0.167 1.433 6.540 0.234 0.974 0.997 0.999 0.054 0.160 2.178 0.081
 • NeRFm. Interpolation 0.739 0.936 0.979 0.168 1.420 6.418 0.231 0.973 0.997 0.999 0.053 0.158 2.178 0.081
 • NeRFm. Angled ±$\pm 3^{\circ}$ 0.721 0.933 0.977 0.171 1.418 6.569 0.238 0.973 0.997 0.999 0.054 0.158 2.165 0.081
 • NeRFm. Horiz. ±$\pm 30\,cm$ 0.730 0.929 0.976 0.168 1.451 6.740 0.238 0.974 0.997 1.000 0.054 0.158 2.158 0.081
 • NeRFm. Vert. ±$\pm 30\,cm$ 0.764 0.937 0.979 0.160 1.355 6.365 0.222 0.974 0.997 1.000 0.053 0.156 2.145 0.080
BinsFormer  • NeRFm. Reconstruction 0.744 0.933 0.977 0.171 1.447 6.476 0.230 0.973 0.997 0.999 0.055 0.159 2.181 0.082
 • NeRFm. Interpolation 0.727 0.933 0.977 0.174 1.462 6.481 0.236 0.973 0.997 0.999 0.054 0.159 2.191 0.081
 • NeRFm. Angled ±$\pm 3^{\circ}$ 0.764 0.940 0.979 0.166 1.395 6.228 0.224 0.975 0.997 0.999 0.054 0.154 2.131 0.080
 • NeRFm. Horiz. ±$\pm 30\,cm$ 0.736 0.933 0.978 0.177 1.501 6.483 0.234 0.973 0.997 0.999 0.054 0.160 2.173 0.081
 • NeRFm. Vert. ±$\pm 30\,cm$ 0.770 0.938 0.978 0.166 1.416 6.181 0.220 0.972 0.997 0.999 0.055 0.161 2.187 0.082
4.5 Training of monocular depth estimation models
To demonstrate that our augmentation technique is model-agnostic, we show it in combination with three popular MDE architectures: AdaBins [5], DepthFormer [17], and BinsFormer [18]. We use the Monocular-Depth-Estimation-Toolbox [16], an open-source monocular depth estimation toolbox aimed at collecting several state-of-the-art MDE models into a single environment. We train the described off-the-shelf MDE networks on the augmented KITTI Eigen [12, 8] train split and compare the results to Classic, FDA, and Style Transfer augmentation results. Due to computational constraints, we cannot use the original batch size used for DepthFormer and BinsFormer. Hence, following the linear scaling rule described in [13], we decrease the learning rate and increase the number of iterations accordingly. Other hyperparameters are kept as described in the corresponding papers to show the influence of the augmentation methods. For the evaluation of the MDE models, we use the same metrics as in [3]. For their definitions, we refer to the supplementary material.
@@ 4.5 Training of monocular depth estimation models
4.6 Evaluation of zero-shot dataset transfer to Waymo
@@
Evaluation on a different dataset provides a good insight into the models’ generalization capabilities. The models are strictly trained on KITTI and KITTI-based NeRFmented images in this comparison. Tab. 1 shows that the NeRFmented runs deliver massive uplifts during zero-shot evaluation on the Waymo dataset, beating baseline performance in a large majority of metrics for every single augmentation strategy. Note that the NeRFmented results shown in Tab. 1 correspond to the augmentation regime (iv), vertical translation. We furthermore note that NeRFmentation performs better when classic augmentations are turned off.
RGB
Refer to caption Refer to caption
DepthFormer
Refer to caption Refer to caption
NeRFmented DepthFormer
Refer to caption Refer to caption
Figure 6: Qualitative results on the KITTI [12] dataset. We show the qualitative depth predictions of the vanilla-trained DepthFormer [17] vs our proposed NeRFmented DepthFormer, demonstrating the capability of our method to recover fine-grain details in the prediction that the baseline is not able to predict. Color scale: 0 (purple) to 80 meters (yellow).
As shown in Tab. 2, out of all strategies, the vertical translation strategy performs best. We believe that it forces the MDE to overfit less to a specific camera height, allowing better performance for the different car heights in Waymo. Generally, the largest difference in performance lies between NeRFmented and non-NeRFmented runs, whereas it doesn’t seem to matter as much which NeRFmentation strategy is picked. We hypothesize that since all of the NeRFmentation runs offer depth completion, it is this additional dense supervision that helps these models perform. However, we also note that in most cases vertical translation outperforms reconstructing the RGB-D pairs. Therefore, we argue that it is not only the depth completion offered by NeRFs that increases model robustness but also the additional viewpoints. The augmentation’s quantitative success is mirrored by the qualitative results displayed in Fig. 5. An increase in fine detail prediction in street signs, stoplights, streetlamps, and foliage is clearly visible. Car outlines become sharper and background building details are much improved as well.
Table 3: Comparison of data augmentation methods on the KITTI Eigen [12, 8] test split. We showcase the performance of AdaBins [5], DepthFormer [17] and BinsFormer [18] models trained on the KITTI Eigen [12, 8] train split and compare different augmentation methods to our NeRFmentation method. CAug indicates whether classic augmentations were used during training. For NeRFmentation, the NeRFm. Vert. ±$\pm 30cm$ strategy is shown. The best results are in bold. The second best results are underlined, and the baseline is highlighted.
CAug. ✓ Adabins [5] 0.964 0.995 0.999 0.058 0.190 2.360 0.088
 • FDA [39] 0.962 0.995 0.999 0.059 0.200 2.385 0.090
 • Style Transfer [1] 0.960 0.994 0.999 0.060 0.211 2.482 0.093
 • NeRFmentation 0.960 0.994 0.999 0.062 0.210 2.427 0.094
CAug. ✗ Adabins [5] 0.962 0.995 0.999 0.060 0.204 2.449 0.092
 • FDA [39] 0.956 0.994 0.999 0.062 0.219 2.542 0.096
 • Style Transfer [1] 0.948 0.992 0.998 0.066 0.253 2.798 0.103
 • NeRFmentation 0.958 0.994 0.999 0.062 0.215 2.534 0.095
CAug. ✓ DepthFormer [17] 0.975 0.997 0.999 0.052 0.158 2.143 0.079
 • FDA [39] 0.973 0.997 0.999 0.052 0.154 2.123 0.079
 • Style Transfer [1] 0.972 0.997 0.999 0.052 0.160 2.192 0.082
 • NeRFmentation 0.976 0.997 0.999 0.052 0.150 2.088 0.078
CAug. ✗ DepthFormer [17] 0.972 0.997 0.999 0.054 0.159 2.172 0.082
 • FDA [39] 0.971 0.997 0.999 0.054 0.164 2.213 0.082
 • Style Transfer [1] 0.965 0.995 0.999 0.057 0.194 2.525 0.089
 • NeRFmentation 0.974 0.997 1.000 0.053 0.156 2.145 0.080
CAug. ✓ BinsFormer [18] 0.974 0.997 0.999 0.052 0.151 2.098 0.079
 • FDA [39] 0.974 0.997 0.999 0.053 0.155 2.117 0.080
 • Style Transfer [1] 0.971 0.997 0.999 0.054 0.165 2.226 0.083
 • NeRFmentation 0.973 0.997 0.999 0.054 0.156 2.124 0.081
CAug. ✗ BinsFormer [18] 0.973 0.997 0.999 0.054 0.159 2.191 0.081
 • FDA [39] 0.973 0.997 0.999 0.054 0.159 2.178 0.081
 • Style Transfer [1] 0.968 0.996 0.999 0.058 0.184 2.411 0.088
 • NeRFmentation 0.972 0.997 0.999 0.055 0.161 2.187 0.082
4.7 Evaluation on KITTI
When evaluating NeRFmentation on the KITTI Eigen [12, 8] test split, large qualitative improvements can be seen (Fig. 6). As a result of the densification of the depth maps, the MDE architecture gains much-needed supervision here. This effect is apparent by the increased clarity in the upper parts of the predicted depth maps. We also notice increased detail in the entire image for objects such as street signs, poles, and trees.
The quantitative results on the KITTI Eigen test split are displayed in Tab. 3. For all NeRFmentation results shown, we choose the vertical translation strategy for consistency. NeRFmented models perform slightly worse than baseline on average. Better small detail prediction or improved performance in the upper part of the image are not measured as the LiDAR-generated depth maps do not offer any ground truth information in the upper parts of the image (see Fig. 4). Therefore, areas, where NeRFmentation provides denser depth information, are not captured by the evaluation metrics on this dataset. To eliminate this factor in evaluation, we also evaluate our models on our own densified (NeRFmented) KITTI test set. Here, Tab. 4 of the main paper shows NeRFmentation significantly outperforming the baseline. In Tab. 2 we notice that the same NeRFmentation strategies which work well on Waymo continue to do so for KITTI. We also note that contrary to the case when evaluated on Waymo, classic augmentations increase the performance when using NeRFmentation. Interestingly some models benefit more from the additional viewpoints and dense depth information, while others seem to be more sensitive to the slight domain shift introduced by the synthetic data.
Table 4: Comparison of performances on the NeRF-generated KITTI dataset. This table shows the performance of baseline and NeRFmented KITTI trained models evaluated on a NeRF rendered subset of the KITTI Eigen [8, 12] test split scenes. NeRFmentation improves upon baseline performance independent of the architecture used across all strategies. No classic augmentation is used. The best results are in bold. The second best results are underlined, and the baseline is highlighted.
Augmentation $\delta_{1}\uparrow$ $\delta_{2}\uparrow$ $\delta_{3}\uparrow$ REL $\downarrow$ SQ. REL $\downarrow$ RMS $\downarrow$ RMSLOG $\downarrow$
Adabins [5]  • No Augm. 0.776 0.964 0.989 0.182 1.011 4.104 0.205
 • Classic 0.761 0.961 0.988 0.183 0.987 4.248 0.211
 • NeRFm. Reconstruction 0.909 0.986 0.995 0.126 0.639 3.380 0.137
 • NeRFm. Interpolation 0.908 0.986 0.994 0.121 0.615 3.377 0.137
 • NeRFm. Angled ±$\pm 3^{\circ}$ 0.898 0.985 0.994 0.140 0.821 3.488 0.142
 • NeRFm. Horiz. ±$\pm 30cm$ 0.902 0.985 0.995 0.140 0.823 3.454 0.141
 • NeRFm. Vert. ±$\pm 30cm$ 0.850 0.984 0.995 0.170 0.852 3.646 0.172
DepthFormer [17]  • No Augm. 0.801 0.977 0.992 0.168 0.805 3.783 0.189
 • Classic 0.791 0.978 0.992 0.171 0.840 3.804 0.193
 • NeRFm. Reconstruction 0.922 0.990 0.997 0.109 0.492 3.073 0.125
 • NeRFm. Interpolation 0.923 0.989 0.996 0.104 0.455 3.073 0.125
 • NeRFm. Angled ±$\pm 3^{\circ}$ 0.919 0.989 0.996 0.112 0.521 3.119 0.127
 • NeRFm. Horiz. ±$\pm 30cm$ 0.922 0.989 0.996 0.114 0.532 3.186 0.128
 • NeRFm. Vert. ±$\pm 30cm$ 0.905 0.990 0.996 0.143 0.583 3.350 0.153
BinsFormer [18]  • No Augm. 0.799 0.979 0.992 0.172 0.846 3.744 0.190
 • Classic 0.807 0.979 0.992 0.168 0.831 3.743 0.189
 • NeRFm. Reconstruction 0.920 0.989 0.996 0.109 0.476 3.066 0.127
 • NeRFm. Interpolation 0.925 0.989 0.996 0.106 0.454 3.010 0.124
 • NeRFm. Angled ±$\pm 3^{\circ}$ 0.921 0.988 0.996 0.113 0.514 3.061 0.127
 • NeRFm. Horiz. ±$\pm 30cm$ 0.926 0.990 0.996 0.111 0.509 3.101 0.125
 • NeRFm. Vert. ±$\pm 30cm$ 0.897 0.988 0.996 0.145 0.585 3.334 0.158